Determine the capacitance of a capacitor whose stored charge is 300uC and is supplied with a 20V battery.

a. 6000F

b. 15F

c. 6x10^-3F

d. 1.5x10^-3F
Answer #1

We know, stored charge Q=CV, where C is the capacitance of the capacitor and V is the supplied voltage.

Here, C=300 microCoulomb= 300×10-6 C

And V= 20V

Therefore, C 300 * 10-6 -=1.5 * 10-5F =